A rectangular bar of carbon is 1cm by 1cm cross sectional area and 50cm long.

I. Find the resistance between it's square end.

II. The resistance between two opposite rectangular phase. Resistivity= 3.5x10^-5ohms


1
Expert's answer
2022-02-24T11:30:53-0500

Answer


resistance is given by

"R=\frac{\rho l}{A}\\=\frac{3.5*10^{-5}*0.50}{0.0001}\\=0.175\Omega"